OpenStudy (lamechildidc):

Can somebody please help me :/ Simplify (8j^3 – 10j^2 – 7) – (6j^3 – 10j^2 – j + 12). (I got the first one but I'm not sure) 2j^3 – 20j^2 – j + 5 2j^3 + j – 19 2j^3 – 20j^2 + j – 19 2j^3 – j + 5

OpenStudy (lamechildidc):

@Preetha

OpenStudy (lamechildidc):

@Michele_Laino

OpenStudy (michele_laino):

here we have to sum similar terms, and we have to apply the distributive property of multiplication over addition , in order to cancel the parentheses, so we can rewrite your expression as below: \( (8j3 – 10j2 – 7) – (6j3 – 10j2 – j + 12)= 8j3-6j3 -10j2+10j2+j -7-12=...\)

OpenStudy (lamechildidc):

Okay i understand that and then i combined like terms and got the first answer

OpenStudy (michele_laino):

please, first combine like terms, and write what you get

OpenStudy (lamechildidc):

2j3-20j2+j+5

OpenStudy (michele_laino):

we have: \(-10j^2+10j^2=0\)

OpenStudy (michele_laino):

and \(-7-12=-19\)

OpenStudy (lamechildidc):

So it would be B then?

OpenStudy (michele_laino):

correct!

OpenStudy (lamechildidc):

Okay thank you I just got really confused because I know you have to switch the signs but idk lols thanks :)

OpenStudy (michele_laino):

:)